OpenStudy (anonymous):

The volume of a right circular cylinder is 12π in cubed. express the height as a function of the radius. Express the total surface area as a function of the radius.

OpenStudy (anonymous):

Ok, so we know \[ \text{cylinder_height} = h \\ v = \text{base_area} \cdot h = 12 \pi \\ \text{base_area} = \pi \cdot r^2 \\ f(r) =h = \frac{ \text{base_area} \cdot h }{\text{base_area}} = \frac{v}{\text{base_area}} = \frac{12\pi}{\pi \cdot r^2} = \frac{12}{r^2} \\ g(r) = \text{surface} = \text{base_circumference} \cdot h = 2 \pi r \cdot \frac{12}{r^2} = \frac{24 \pi}{r} \]